# Universal Scaling Solution for a Rigidity Transition: Renormalization Group Flows Near the Upper Critical Dimension

### v1

@@ Appendix C Scaling variables in terms of isotropic elastic sheet parameters
Here we expound upon the calculation mentioned in the main text that eventually leads to the universal scaling functions for the vanishing modulus. First, we write the dynamical matrix and Green’s function for an isotropic elastic sheet, with Lamé parameters $\lambda$ and $\mu$. This is typically done in the reverse order using the constitutive relations (i.e. coarse-graining a lattice system with dynamical matrix $\mathbf{D}$), recovering the long-wavelength stiffness tensor $K_{ijk\ell}$. For a lattice with real-space unit cell volume $V$ and $n$ sites in the unit cell, we have
@@ C.5 Density of states scaling
This remarkably simple scaling form is true within the CPA, but not in general. Setting $G=0$ allows us to exactly evaluate this; in this scaling limit there is only a nonzero density of states for $\Omega>1/2$ and it is flat at high frequency. We have
@@
Figure 5: Density of states for a nearly floppy 3D viscoelastic membrane. The dangerously irrelevant variable controls the low-frequency density of states in the undamped case. See Figure 2 for the excess density of states $\mathcal{D}\left(\Omega,G\right)/\Omega^{2}$.
@@ C.6 Green’s function scaling
These forms imply diverging length scales at the transition: $\ell^{*}\sim\left|\delta p\right|^{-1/2}$ for $d>2$, $\ell^{*}\sim\left|\delta p\right|^{-1/d}$ for $d<2$, and $\ell^{*}\sim\left|\delta p\right|^{-1/2}\left|\log\left|\delta p\right|\right|^{1/2}$ in $d=2$ (noted also in Table 2). If a phonon has a wavelength shorter than $\ell^{*}$, it is strongly overdamped. In the case of jamming, the transverse shear mode is associated with a diverging length scale of the type analyzed here and elsewhere [16, 38] but the longitudinal mode has a different scaling.
@@ Appendix D Details of the triangular lattice numerics
Here we directly compare our continuum, isotropic expansion of the CPA to the lattice CPA for the bond-diluted triangular lattice. To make this comparison with no numerically determined fitting parameters, we must know the values of the nonuniversal constants $\mu_{0}$, $f_{02}\equiv\lim_{d\rightarrow 2^{+}}f_{0}\Gamma\left(d/2-1\right)$, and $g_{2}/g_{02}$. We take the triangular lattice with nearest-neighbor bonds of strength $k$ and bond length $a=1$. The dynamical matrix is

### v2

@@ Appendix C Scaling variables in terms of isotropic elastic sheet parameters
Here we expound upon the calculation mentioned in the main text that eventually leads to the universal scaling functions for the vanishing modulus. We take the perspective of a long-wavelength limit of a bond percolation CPA that is explicitly isotropic at long-wavelengths; this recovers the CPA for amorphous spring networks developed in [17]. First, we write the dynamical matrix and Green’s function for an isotropic elastic sheet, with Lamé parameters $\lambda$ and $\mu$. This is typically done in the reverse order using the constitutive relations (i.e. coarse-graining a lattice system with dynamical matrix $\mathbf{D}$), recovering the long-wavelength stiffness tensor $K_{ijk\ell}$. For a lattice with real-space unit cell volume $V$ and $n$ sites in the unit cell, we have
@@ C.5 Density of states scaling
This remarkably simple scaling form is true within the CPA [48], but not in general. Setting $G=0$ allows us to exactly evaluate this; in this scaling limit there is only a nonzero density of states for $\Omega>1/2$ and it is flat at high frequency. We have
@@
Figure 5: Density of states for a nearly floppy 3D viscoelastic membrane. The dangerously irrelevant variable controls the low-frequency density of states in the undamped case. The rescaled density of states has a square-root cusp at $G=0$ $(\delta p=0)$ [12, 16]; as discussed in section III.1, the continuum CPA, and our universal scaling limit, do not show the $\omega^{4}$ contribution to the density of states found from quasilocalized modes [49]. See Figure 2 for the excess density of states $\mathcal{D}\left(\Omega,G\right)/\Omega^{2}$.
@@ C.6 Green’s function scaling
These forms imply diverging length scales at the transition: $\ell_{c}\sim\left|\delta p\right|^{-1/2}$ for $d>2$, $\ell_{c}\sim\left|\delta p\right|^{-1/d}$ for $d<2$, and $\ell_{c}\sim\left|\delta p\right|^{-1/2}\left|\log\left|\delta p\right|\right|^{1/2}$ in $d=2$ (noted also in Table 2). If a phonon has a wavelength shorter than $\ell_{c}$, it is strongly overdamped. In the case of jamming, the transverse shear mode is associated with a diverging length scale of the type analyzed here and elsewhere [16, 39] but the longitudinal mode has a different scaling.
@@ Appendix D Details of the triangular lattice numerics
Here we directly compare our continuum, isotropic expansion of the CPA to the lattice CPA for the bond-diluted triangular lattice. (As noted earlier, the rigidity transition of the diluted triangular lattice is not described correctly by our CPA analysis [22]. The static critical exponents for the triangular lattice have lengths which scale as $\left|\delta p\right|^{-\nu}$ and moduli which scale as $\left|\delta p\right|^{f}$, with $\nu\sim 1.3\pm 0.2$ and $f\sim 2.2\pm 0.3$; CPA predicts $\nu=1/2$ and $f=1$, and log corrections.)
To make this comparison with no numerically determined fitting parameters, we must know the values of the nonuniversal constants $\mu_{0}$, $f_{02}\equiv\lim_{d\rightarrow 2^{+}}f_{0}\Gamma\left(d/2-1\right)$, and $g_{2}/g_{02}$. We take the triangular lattice with nearest-neighbor bonds of strength $k$ and bond length $a=1$. The dynamical matrix is